Consider the dataset below for lead residue as an effect of number of pistons in an engine:

Number of Pistons Lead (Parts per Million)

2 (45)

4 (38)

6 (24)

8 (32)

10 (12)

12 (4)

Using SPSS, create a scatterplot of the data, graph the least-squares regression line, and find the correlation. Are there any influential points (and how do you recognize them)? How many parts per million of lead should we expect when there are five pistons? Be sure to export your outputs to a Word document in order to edit the graphs and statistics together and remove unnecessary computer code and notes.
Answer #1
SUMMARY OUTPUT
Regression Statistics
Multiple R 0.9392
R Square 0.8820
Adjusted R Square 0.8526
Standard Error 6.0099
Observations 6
ANOVA
df SS MS F Significance F
Regression 1 1080.3571 1080.3571 29.9110 0.0054
Residual 4 144.4762 36.1190
Total 5 1224.8333
Coefficients Standard Error t Stat P-value Lower 95%
Intercept 53.3333 5.5949 9.5325 0.0007 37.7993
x -3.9286 0.7183 -5.4691 0.0054 -5.9230

y^ = 53.3333 -3.9286 *x

r = -0.9392

when x= 5

y^ = 53.3333 -3.9286 *5 = 33.6903